OpenStudy (anonymous):

the top of a ladder leaning against a building reaches a point on the building which is 29 feet above the ground. if the base of the ladder is 7 feet from the building, what is the measure of the angle that the ladder makes with the level ground to the nearest degree

OpenStudy (campbell_st):

start by drawing a diagram  which sides of the triangle have measurements... opposite, adjacent, hypotenuse?

OpenStudy (anonymous):

opposite & adjacent

OpenStudy (campbell_st):

great... so which trig ratio uses opposite and adjacent..?

OpenStudy (anonymous):

tan

OpenStudy (campbell_st):

great so you can say \[\tan(x) = \frac{29}{7}\] can you find the angle from here..?

OpenStudy (anonymous):

would i divide 29 by 7 and get that number and then plug it into x for tan?

OpenStudy (anonymous):

the answer i got is 0.07243277332 which i don't think is correct

OpenStudy (campbell_st):

not quite to find and angle, you need to use the inverse so if \[\tan(x) = \frac{29}{7} ~~~then~~~~ x = \tan^{-1}(\frac{29}{7})\]

OpenStudy (anonymous):

i am still confused on how i would solve it though

OpenStudy (campbell_st):

ok... does your cacluator have a shift or 2nd key..?

OpenStudy (anonymous):

2nd key

OpenStudy (campbell_st):

ok... press 2nd and then tan the screen would show \[\tan^{-1}(\] is that correct..?

OpenStudy (campbell_st):

or ir may show atan(

OpenStudy (campbell_st):

did you get either of those on the display..?

OpenStudy (anonymous):

sorry my computer died let me try now

OpenStudy (anonymous):

it shows atan(

OpenStudy (campbell_st):

ok.... that's fine, that is how you find the angle measure.... so of you press clear calculator the to find the angle you need 2nd tan then inside the open bracket enter 29/7 and close the brackets... then press = or enter... which ever you may have

OpenStudy (anonymous):

i got 79.42956...

OpenStudy (campbell_st):

umm...can you check you're calculate has a small D for degrees in the display... its just I got 76.42..

OpenStudy (campbell_st):

or did you make a typo

OpenStudy (anonymous):

no i made a type

OpenStudy (campbell_st):

ok... so that looks good... atan, asin and acos are used to find an angle... just put the fraction in the brackets

OpenStudy (anonymous):

okay thank you so much